### v1

# Now at ]Western Digital Research Center, Western Digital Corporation, San Jose, CA 95119, USA

Abstract
The topological semimetal CrPt3 has potential for generating unconventional spin-torques due to its ferrimagnetic ordering, topological band structure, and high anomalous Hall effect. CrPt3 exhibits ferrimagnetic behavior only in its chemically ordered phase and is paramagnetic in its chemically disordered phase. By controlling the growth and annealing temperatures epitaxial films of both chemically ordered and disordered phases of CrPt3 are prepared allowing us to investigate the role of magnetic ordering on unconventional torque generation. We use angle dependent spin-torque ferromagnetic resonance to probe the spin-torques generated from epitaxial CrPt3 in CrPt3/Cu/Ni81Fe19 heterostructures. With current applied along specific directions with respect to the magnetic and crystal order we reveal unconventional spin torques in both ordered and disordered films. When current flows parallel to the $[1\overline{1}1]$ and $[\overline{1}11]$ directions we observe an unconventional field-like torque that is opposite in sign for the two directions, which lack a mirror plane thus allowing unconventional torques to be generated.
I Introduction
Spin-orbit torques play a crucial role in the development of new technology toward the efficient control of magnetization states using charge currents [1]. The spin Hall effect is an efficient way of converting a charge current to a spin current and creating spin-torques on an adjacent ferromagnetic layer. These spin-torques are key to novel information technologies, ranging from power efficient and ultrafast nonvolatile memories[2] to new computational paradigms based on magnetic oscillators emulating neuromorphic functionalities[3].
In the conventional spin Hall effect induced spin-orbit torque geometry, a charge current along the in-plane $x$-direction generates a spin-current along the out-of-plane $z$-direction with a spin polarization along the in-plane $y$-direction ($\bm{\sigma_{y}}$) perpendicular to both charge and spin current. The torques that are generated from this spin-polarization are in-plane damping-like torques of the form $\bm{\tau_{ip}}\propto\textbf{m}\times(\textbf{m}\times\bm{\sigma_{y}})$ and out-of-plane field-like torques of the form $\bm{\tau_{oop}}\propto\textbf{m}\times\bm{\sigma_{y}}$, with m being the magnetization. These torques can only field-free and deterministically switch in-plane magnetization parallel to the spin-polarization direction, i.e., the $y$-direction. However, in ordinary spin-transfer torque memory devices, it is well known that in order to have low switching currents while maintaining high thermal stability a perpendicular magnetic anisotropy, i.e., along the $z$-direction, is preferred, since this avoids demagnetizing fields during the precessional switching[4].
In recent years, interest in generating spin-currents with a spin-polarization that varies from the conventional geometry has emerged, such that other magnetization directions can be switched deterministically. It has been shown that by symmetry breaking through crystal or magnetic structure, unconventional orientations of spin-polarizations are allowed, such that unconventional torque geometries are possible. In WTe2 the reduced crystal symmetry leads to an out-of-plane spin polarization [5], whereas in Mn3Ir and Mn3GaN the non-collinear antiferromagnetic order reduces the symmetry to allow unconventional spin-torques [6, 7, 8]. In epitaxial CoPt/CuPt heterostructures, the reduced symmetry at the interface can also result in unconventional spin-torques [9]. These unconventional spin-torques enable field-free deterministic switching of memory devices with perpendicular anisotropies[9], and may also enable novel magnetization dynamics, such as multiple dynamic droplets[10].
Another material that shows potential for generating such torques is the topological semimetal CrPt3 due to its magnetic ordering [11]. CrPt3 exhibits ferrimagnetic behavior in its chemically ordered phase and paramagnetic behavior in its chemically disordered phase [12]. In the ferrimagnetic phase, the magnetic ordering may give rise to unconventional torques by generating a spin polarization with a polarization component collinear to the magnetization [13, 14, 15, 16, 17].
In this work, we use angle dependent spin-torque ferromagnetic-resonance [18] measurements with current applied along varying crystallographic directions and in both ferrimagnetic and paramagnetic phases to investigate the role of magnetic ordering on unconventional torque generation. We show an unconventional field-like spin torque originating from a spin polarization component along the applied current direction in both chemically ordered and disordered CrPt3 thin films. The presence of unconventional torques in both phases suggests that the crystal structure is the origin of these torques rather than the magnetic ordering.
II Methods
II.1 Sample growth and characterization
CrPt3 was grown on single-crystal MgO (220) substrates using a dc magnetron sputtering system with a base pressure of 5 $\times$ 10-8 Torr. Cr and Pt targets were co-sputtered using a confocal geometry in an Ar environment with sputtering pressure of 2.7 mTorr. The co-sputtering growth rate was 1.5 Å/s. CrPt3 films were grown at 500C to achieve epitaxial growth and control the surface roughness. To obtain chemical ordering, the samples were post-annealed in situ at 850C for 40 minutes. In a second batch of samples, the post-annealing was skipped to obtain chemically disordered films. The samples were naturally cooled in situ followed by Cu and permalloy (Py, Ni81Fe19, Ni and Fe in wt%) deposition at room temperature to minimize interfacial mixing. Lastly, an Al2O3 capping layer was deposited to prevent oxidation. The Cu spacer layer served to magnetically decouple the CrPt3 and Py layers, while the Py layer was used as a sensing layer for spin-torque measurements. The final sample structures were MgO/CrPt3(15)/Cu(3)/Py(8)/Al2O3(2) for the ordered and MgO/CrPt3(15)/Cu(3)/Py(5)/Al2O3(2) for the disordered sample. The numbers in parenthesis indicate film thicknesses in nm. The thicknesses were chosen based on separate resistivity measurements of CrPt3 and Py patterned full films, and normalized resistivity measurements of Cu(3) on patterned Cu/Ge multilayers.
The chemical ordering of CrPt3 was confirmed using out-of-plane X-ray diffraction as shown in Fig. 1(a). While both films show a (220) peak, chemically ordered CrPt3 exhibits a characteristic (110) peak whereas disordered CrPt3 does not. As shown in Fig. 1(c), the full width at half maximum of the $\theta$ scan over CrPt3 (110) is 1.843, indicating good crystalline quality. To confirm epitaxy, in-plane $\phi$ scans were performed, which show two-fold symmetry and alignment with the MgO substrate. [Fig. 1(b)]
Magnetization loops were measured using vibrating sample magnetometry with magnetic fields applied along different crystal directions as shown in Fig. 1(d). The measurements show that the magnetic easy axis is along $[\overline{1}10]$ with a coercive field of $\mu_{0}H_{c}=0.3$ T. The (110) growth direction was specifically chosen, such that the magnetic easy axis of CrPt3 lies in plane and current can be applied along varying in-plane directions with respect to the easy axis to understand the role of magnetism on the spin-torque geometry. The large coercive field allows for the magnetization of CrPt3 to be set using a correspondingly large magnetic field and not be changed during spin-torque ferromagnetic resonance measurements with a smaller magnetic field, where only the Py magnetization is rotated. The maximum magnetic field used during measurements was $\mu_{0}H_{c}=0.135$ T.
Figure 1: (a) Out-of-plane X-ray diffraction measurement of chemically ordered (blue) and disordered (brown) CrPt3, (b) In-plane X-ray diffraction of chemically ordered CrPt3 with 2-fold symmetry aligned with the MgO (220) substrate (c) Rocking curve over CrPt3 (110) diffraction peak, and (d) room temperature vibrating sample magnetometry with magnetic field along the in-plane $[\overline{1}10]$ (blue), $[001]$ (red), and out-of-plane (green) directions.
II.2 Spin torque measurements
[…]
Spin-torque ferromagnetic-resonance measurements were used to characterize the spin-torques generated in CrPt3. In these measurements, an 8 GHz rf current is applied through the device which generates an rf spin current in the CrPt3 that is injected through the Cu into the Py. There, the spins exert a torque on the magnetization of the Py, and together with Oersted fields drive the magnetization into precession around an in-plane applied magnetic field when the magnetic field matches the resonance field of the Py. This leads to an rf resistance modulation due to the anisotropic magnetoresistance in Py. The mixing of the rf resistance and the applied rf current generates a dc mixing voltage, which is measured using a lock-in amplifier.
Figure 2: (a) Crystal and magnetic structure of chemically ordered (right) and disordered (left) CrPt3, where blue spheres represent Cr and gray spheres represent Pt atoms. The gray plane corresponds to the film plane and the red arrow shows the magnetic easy axis direction. (b) Schematic of the device structure used for spin-torque ferromagnetic resonance measurements with $\tau_{oop}$ corresponding to out-of-plane torques manifested in the antisymmetric and $\tau_{ip}$ corresponding to in-plane torques manifested in the symmetric Lorentzian component. $\phi$ is the angle between the rf current ($I_{rf}$) and the magnetization ($M$) that is determined by the magnetic field ($H$). (c) Schematic of the four crystal orientations in which rf current is applied. $\theta$ represents the angle between the $[001]$ direction and the applied current ($I_{rf}$, yellow arrow) and $\phi$ the angle between the applied current and the magnetic field ($H$, green arrow). The red arrow shows the magnetic easy axis direction. (d) Exemplary field sweep at a single angle. The resonance lineshape can be decomposed into antisymmetric and symmetric components.
The magnetic field is swept through the resonance condition of the Py to measure the resonance line shape, which is composed of a symmetric and antisymmetric Lorentzian function and can be written as
$V_{mix}$ $=S\frac{\Delta H^{2}}{(H_{ext}-H_{0})^{2}+\Delta H^{2}}$ (1)
$+A\frac{\Delta H(H_{ext}-H_{0})}{(H_{ext}-H_{0})^{2}+\Delta H^{2}}$
Here, $\Delta H$ is the half width at half maximum of the resonance linewidth, $H_{ext}$ is the applied magnetic field, $H_{0}$ is the Py resonance field, and S and A are the symmetric and antisymmetric Lorentzian components, respectively, that measure the size of the spin-torques. An exemplary fit to a field sweep is shown in Fig. 2(d). The antisymmetric component corresponds to out-of-plane torques, while the symmetric component corresponds to in-plane torques [$\tau_{oop}$ and $\tau_{ip}$ respectively in Fig. 2(b)].
To fully analyze the spin-torque symmetry, the magnetic field is rotated through a 360 in-plane angle to vary the angle between the applied current and Py magnetization [denoted by $\phi$ as shown in Fig. 2(c)], and a magnetic field sweep is performed every 5. The size of the symmetric and antisymmetric components depends on the relative orientation between the spin-polarization and Py magnetization, such that the spin-torque and thereby spin-polarization direction can be extracted from the angular dependence of the voltage signals.
III Results
Figure 3: Antisymmetric Lorentzian component of the resonance signal as a function of the angle between the current and magnetic field for the ordered CrPt3 devices at (a) 0, (b) 90, (c) +54.7, and (d) -54.7 with respect to the $[\overline{1}10]$ crystal direction of CrPt3. The red line shows the fit of the data shown with black circles. For (c) and (d) the blue (green) lines show the contribution from conventional (unconventional) torque components.
Figure 4: Antisymmetric Lorentzian component of the resonance signal as a function of the angle between the current and magnetic field for the disordered CrPt3 devices at (a) 0, (b) 90, (c) +54.7, and (d) -54.7 with respect to the $[\overline{1}10]$ crystal direction of CrPt3. The red line shows the fit of the data shown with black circles. For (c) and (d) the blue (green) lines show the contribution from conventional (unconventional) torque components.
Table 1: Conventional $A_{FL}^{Y}$ and unconventional $A_{FL}^{X}$ fitting parameters in $\mu$V.
Ordered $[001]$ $[\overline{1}10]$ $[\overline{1}11]$ $[1\overline{1}1]$
$A_{FL}^{Y}$ 7.41 5.56 6.04 4.23
$A_{FL}^{X}$ 0 0 0.72 -0.35
$A_{FL}^{X}$/$A_{FL}^{Y}$ 0 0 0.12 -0.08
Disordered $[001]$ $[\overline{1}10]$ $[\overline{1}11]$ $[1\overline{1}1]$
$A_{FL}^{Y}$ 2.35 2.90 2.28 2.18
$A_{FL}^{X}$ 0 0 0.23 -0.16
$A_{FL}^{X}$/$A_{FL}^{Y}$ 0 0 0.10 -0.07
Figure 3 shows the angle dependent antisymmetric mixing voltage component of the ordered ferrimagnetic CrPt3 sample. The signal from the devices oriented along $[\overline{1}10]$ (90) and $[001]$ (0) [Figs. 3(a) and (b), respectively] can be fitted with $A=A_{FL}^{Y}\mathrm{sin}(2\phi)\mathrm{cos}(\phi)$, where $A_{FL}^{Y}$ is related to the strength of the field-like torque originating from conventional $y$-direction polarized spins and Oersted fields. To fit the signal from the devices oriented along $[\overline{1}11]$ (+54.7) and $[1\overline{1}1]$ (-54.7) an additional $A_{FL}^{X}\mathrm{sin}(2\phi)\mathrm{sin}(\phi)$ component needs to be added, where $A_{FL}^{X}$ is related to the strength of a field-like torque originating from unconventional spins polarized parallel to the current direction. The sign of this component is opposite in the $[\overline{1}11]$ and $[1\overline{1}1]$ devices.
Figure 4 shows the angle dependent antisymmetric mixing voltage for the chemically disordered paramagnetic CrPt3 sample. Here, the $[\overline{1}10]$ and $[001]$ [Figs. 4(a) and (b), respectively] can again be fitted with the conventional $A=A_{FL}^{Y}\mathrm{sin}(2\phi)\mathrm{cos}(\phi)$ term, while the $[\overline{1}11]$ and $[1\overline{1}1]$ devices require an additional $A_{FL}^{X}\mathrm{sin}(2\phi)\mathrm{sin}(\phi)$ term just like the devices in the ordered sample. Table 1 shows a summary of the fitting parameters used for each sample and device.
If the unconventional torque originated from the magnetic ordering of the CrPt3 we would expect the spin polarization to have a component along the magnetization direction. This would result in a purely conventional spin polarization in the device with current along the $[001]$ direction since the magnetization of the CrPt3 is perpendicular to the current direction, which is what we observed in our measurements. However, with current along the $[\overline{1}10]$ direction, the spin polarization should have an unconventional component along the current direction and collinear to the magnetization. This is not the case for our measurements where we only observe an unconventional spin polarization component with current along the $[\overline{1}11]$ and $[1\overline{1}1]$ directions. This alone suggests that the origin of the unconventional torque is not the magnetic ordering of the CrPt3. Additionally, since we observe an unconventional torque component even in disordered paramagnetic CrPt3, where the magnetic ordering is absent, the unconventional torque cannot come from the magnetic ordering.
Instead, it originates from the broken mirror symmetry due to the crystal structure of CrPt3. When current is applied along the $[1\overline{1}1]$ and $[\overline{1}11]$ directions it flows parallel to a low symmetry axis with a broken mirror symmetry [see Fig. 2(c)]. In contrast, when current is applied along the $[001]$ and $[\overline{1}10]$ directions it flows parallel to a high symmetry axis with the mirror symmetry preserved. The broken symmetry lifts the restriction that charge current, spin current, and spin polarization must be mutually perpendicular and allows unconventional spin polarization directions to exist [19]. Unconventional torques from a broken mirror symmetry due to the crystal structure have been observed in WTe2 [5], MnPd3 [20], and CoPt/CuPt [9], where an unconventional spin polarization is only generated with a current along a low symmetry axis. These results indicate that unconventional spin-orbit torques can be generated even in high symmetry materials as long as the current flow is along a low symmetry crystalline orientation.
IV Conclusion
In summary, we have shown that an unconventional in-plane field-like torque is generated in both ordered (ferrimagnetic) and disordered (paramagnetic) CrPt3. Our spin-torque ferromagnetic resonance measurements reveal that only when current is applied parallel to axes where mirror symmetry is broken ($[1\overline{1}1]$ and $[\overline{1}11]$) is an unconventional torque generated. With current applied parallel to high symmetry axes ($[001]$ and $[\overline{1}10]$) only conventional torques are observed.
V Acknowledgement
This work was supported as part of Quantum Materials for Energy Efficient Neuromorphic Computing (Q-MEEN-C), an Energy Frontier Research Center funded by the US Department of Energy, Office of Science, Basic Energy Sciences under Award No. DE-SC0019273.
VI Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.